OpenStudy (anonymous):

Prove that (14^n + 4)/2 is never prime.

OpenStudy (anonymous):

* (14^n + 11)/2

OpenStudy (mr.math):

The expression you gave is never integer.

OpenStudy (mr.math):

\(14^n+11\) is odd \(\forall n\in \mathbb{Z}\), which implies \(\frac{14^n+11}{2}\) is never integer and hence never prime.

OpenStudy (anonymous):

what about n=0?

OpenStudy (anonymous):

14^n, for n>0 will always end with a 4 or a 6. And since 4+11 = 15 which is odd and 6+11 = 17 which is odd. 14^n + 11 will always be odd. And therefore as Mr.Math said, it won't be an integer and hence not a prime.

OpenStudy (anonymous):

I never thought of that!

OpenStudy (mr.math):

sorry, I was just about to change tjat\(\forall n\in \mathbb{N}\).

OpenStudy (mr.math):

It's obviously not prime for n=0, because (1+11)/2=6.

OpenStudy (anonymous):

Simply even+odd is odd hence the proof.